A firm has a current capital structure consisting of $400,000 of 6 percent annual interest debt...

  1. A firm has a current capital structure consisting of $400,000 of 6 percent annual interest debt and 50,000 shares of common stock. The firm's tax rate is 21 percent on ordinary income. If the EBIT is expected to be $200,000, the firm's earnings per share will be ________.

    $0.74

    $2.78

    $3.12

    $3.52
Answer #1
Calculations Particulars Amount $
A Earnings Before Interest and Taxes ( EBIT )        2,00,000
B = $400,000 x 6% Interest           24,000
C = A - B Earnings Before Taxes ( EBT )        1,76,000
D = C x 21% Taxes           36,960
E = C - D Earnings After Taxes        1,39,040
F Number of shares of common stock           50,000
G = E / F Earnings Per Share ( EPS )               2.78

So, as per above calculations, option B is the correct option